### v1

@@ 4.2 All Deformations
The second most popular deformation class concerns models accounting for all the primary deformations of continuum and soft robots. Below, we review the literature on rod models exhibiting all strains in cable-driven, pneumatic, parallel continuum, and biomechanical structures. Table 2 summarizes the class with the notable works.
4.2.1 Deformations by Cables
Renda et al. (2012) present a geometrically exact static Cosserat model of a conical tendon-driven CSM inspired by the octopus. Cable tension is modeled as a pure tension element, neglecting friction between the cable and the silicone body. After a FEM validation of the algebraic equations, the Cosserat model is validated experimentally for a static planar motion. The following work investigated the dynamic interaction of the silicone arm with a dense medium and the coupled tendon condition, addressing the external forces such as gravity, buoyancy, drag, added mass, and cable load (Renda et al., 2014). The dynamic experimental validation concerned the replication of characteristic octopus movements. In addition, an ablation study suggests that minimal internal viscosity guarantees solution stability and minimizes computational costs. Then, Renda et al. (2018b) introduced a unified dynamic Cosserat rod framework for underwater soft robots, which considers the nonlinearities of soft bodies, offering generality that extends to shells. The framework, which could aid the evaluation of potential soft robot designs, was validated experimentally using an autonomous vehicle with four cable-driven octopus-inspired CSM. At the same time, Renda et al. (2018a) presented an approach called piecewise constant strain (PCS) that discretizes continuous Cosserat rods in pieces, assuming constant strain for each piece. The PCS generalizes the PCC (Webster III and Jones, 2010) providing (i) numerical robustness given by the exponential formulation, (ii) inclusion of shear and torsion due to the CRT, and (iii) possible extension to rigid kinematic chains. The PCS is validated in real-world experiments with a cable-driven siliconic CSM.
Concerning discretization approaches, it was proposed a real-time discretization technique to solve the PDE of a Cosserat rod (Till et al., 2019). The method discretizes the time derivatives and solves the ODEs for the boundary value problem (BVP) in at each timestep. The authors validate the efficient method by comparing the simulation to high-speed footage of a cantilever rod.
For MIS applications, (Ibister et al., 2021) presents a dynamic CRT model for a continuum robot composed of a backbone and disks to support actuating tendons. The model includes the gravity (7) and tip wrenches produced by the tendons. The comparison between simulation and experiments shows a mean tip position error.
Recently, a cable-driven continuum robotic arm was integrated within an aerial manipulation system (Samadikhoshkho et al., 2020). The arm dynamics are modeled with a Cosserat rod with a floating base , incorporating both the actuation and quadrotor effects. They proposed a robust adaptive position controller and proved its stability using the direct Lyapunov stability theorem. The scheme is validated in simulation for quasi-static, passive, and active arm conditions.
4.2.2 Deformations by Pneumatic Actuators
Trivedi et al. (2007, 2008b) present the first dynamic Cosserat rod model for pneumatic CSM based on a fiber-reinforced model of the air muscle actuators. The model includes the effect of material nonlinearities, distributed forces, and payloads and is geometrically exact for large curvature, shear, torsion, and extension. The model is experimentally validated statically on the OctArm V for planar movements with vertical and horizontal base orientation, achieving results about ten times more accurate than the constant curvature model. Subsequently, the model is used within a model-based design optimization study for pneumatic and hydraulic CSM (Trivedi et al., 2008a). Recently, the PCS discretization was combined with Screw Theory, providing a new model of actuation sources for pneumatic actuators (Renda et al., 2017). This approach enables the routing of fluidic actuators (even tendons) along any path in the soft arm. They define a soft geometric Jacobian for the PCS formulation and a technique to identify the active DoFs, applying it to the study of manipulable and force-closure conditions. Finally, they validate the theory on a STIFF-FLOP arm (Cianchetti et al., 2013). Additionally, Roshanfar et al. (2021) described a pneumatic soft robot with a Cosserat rod model, which was solved as an initial value problem with homogeneous Neumann and Dirichlet boundary conditions for a given tip force. The experimental validation measured a tip position error of .
Analytical solutions and detailed FEM simulations are useful means for comparison and validation. For example, analytic solutions for the statics of a cantilever quasi-static Cosserat rod under external loads are derived (Grazioso et al., 2018). The position and orientation fields of the arm are defined on a Lie group, while the velocity and deformation fields are obtained from the time and space derivatives, respectively. A constitutive law between internal forces and strains enables the definition of the arm’s internal strain energy. Therefore, the principle of virtual work provides the static equilibrium equations. Finally, quasi-static rod configurations are compared to analytical solutions for pure bending and shear loads. Conversely, a comparative study for a Cosserat model and a FEM model of a soft pneumatic actuator was presented (Berthold et al., 2021), providing a preliminary parameter identification process. The FEM model was also validated experimentally. Moreover, an operational point-dependent static Cosserat model is obtained from a detailed volume FEM of a fiber-reinforced pneumatic silicone soft robot (Wiese et al., 2022). FEM simulations determine the actuation-dependent curvature, while the rod stiffness properties are identified by applying external forces to the FEM. The validation against FEM and experimental data, under various conditions with and without loads, show that characterizing a single module is sufficient to predict the behavior of a two-module version of the soft robot.
To tackle large deformations of pneumatic soft robots, Mishra et al. (2022) suggest using an explicit fractional-order Bouc-Wen model that considers material nonlinearities, such as hysteresis. The constitutive law incorporated in a Cosserat rod and tested on a Robotino-XT demonstrated superior accuracy to other models (Wang et al., 2022a).
4.2.3 Deformations with Parallel Manipulators
Till and Rucker (2017) addresses the problem of elastic stability in PCM, introducing a condition to evaluate the elastic stability of PCM solutions based on Kirchhoff and Cosserat rods. A numerical test evaluates this condition and a heuristic stability metric. After verifying that the stability test matches other classical results, they experimentally validate the approach on a six-link PCM. The same authors propose a generalized kinetostatics Cosserat rod model (Black et al., 2017). Model linearization provides the manipulator Jacobian, end-effector compliance, input stiffness, and wrench reflectivity matrices. This enabled the analysis of the PCM manipulability and force-sensing accuracy. Using a similar PCM, they achieve accurate actuation-based force sensing and low orientation error.
Conversely, Armanini et al. (2021) extended the PCS parametrization to soft PCM. They present a mathematical tool for modeling closed-chain geometries composed of soft and rigid elements. The formulation is applied to modeling and analyzing the Fin-Ray finger, especially for studying bending in the opposite direction of the external force. Furthermore, this formulation allowed the development of a novel design of the Fin-Ray finger, finding the optimal number of rigid ribs and their optimal orientation in terms of maximum payload and bending capability.
The elegant work of Lilge and Burgner-Kahrs (2022) present a CRT-based kinetostatic modeling framework for tendon-driven PCM, which enabled to efficiently solve the forward, inverse, and velocity kinetostatic problems. In addition, it facilitates the study of kinematic properties like reachability, singularities, manipulability, and compliance. Experiments on a PCM made of four tendon-driven backbones attached via spherical joints to a platform achieved median pose accuracy of and .
4.2.4 Deformations in Bioinspired Modeling
Cosserat rods can be useful tools for biomechanical studies. With application to flapping flight and passive swimming, Boyer et al. (2017) provide a general unified framework for modeling bio-inspired locomotion robots using soft organs. They model soft parts in two ways: (i) the floating frame approach and (ii) the geometrically exact approach. They implemented the two models in a unified general algorithm based on recursive Newton-Euler. The comparison of the two approaches revealed that the former needs fewer coordinates and stronger deformation approximations. Conversely, the latter is more accurate and manages better the external load. The validation of the unified framework simulated two bio-inspired examples of soft locomotion.
Recently, Zhang et al. (2019) realistically simulates musculoskeletal architectures in PyElastica. They assembled heterogeneous, active, and passive Cosserat rods into dynamic structures that mimic bones, tendons, ligaments, fibers, and muscle connectivity. The simulations involve biological and soft robotic case studies, including the human elbow joint, the design and fabrication of swimming and walking millimeter-long bio-hybrid bots, the synthesis of slithering snakes with intricate musculature layouts, and the replication of feathered bird wings - using up to rods. In these applications, the CRT demonstrated versatility to advance forward and inverse bioengineering designs and fundamental discoveries in the functioning of living organisms.
Boyer et al. (2022) investigates the relationship between CRT and optimal control theory, deriving the statics and dynamics of a continuum robot as solutions to a minimization problem. With this different formulation, the authors revealed significant theoretical insights, such as the intrinsic singularity of the dynamic simulation of continuum robots. Furthermore, the iterative solution of the optimal control problem corresponds to the continuous version of the Newton-Euler algorithm for rigid kinematic chains. Finally, the presented approach is extended to locomotors and applied to a bio-inspired continuum swimmer, validating the theoretical work through simulations.
Subclass | References | Rod Model | Contribution |
Cable | Renda et al. (2014) | CRT | Studied the dynamic interaction of an octopus-inspired CSM with a dense medium. |
Pneumatic | Trivedi et al. (2008b) | CRT | Model validation of CSM with fiber-reinforced actuators varying base orientation. |
Parallel | Till and Rucker (2017) | CRT | Studied and experimentally validated the elastic stability of PCM. |
Armanini et al. (2021) | CRT | Modeled closed-chain soft PCM with validation on a soft-rigid finger. | |
Lilge and Burgner-Kahrs (2022) | CRT | Studied reachability, singularities, manipulability, and compliance of PCM. | |
Bioinspired | Zhang et al. (2019) | CRT | Simulates musculoskeletal systems assembling heterogeneous active/passive rods. |
Boyer et al. (2022) | CRT | Studied the link between CRT and optimal control simulating locomotors and swimmers. |
@@ 4.3 Bend & Twist
Like bending, various robot types exhibit the combination of bending and twisting induced by diverse actuation types (Fig. 5). Table 3 summarizes the Bend & Twist deformation class.
4.3.1 Bend & Twist by Cables
Bending and twisting are usually induced by cables and tendons through specific routing paths (Fig. 5). Rucker and Webster III (2011) present Cosserat rod models for the forward kinematics, statics, and dynamics of a tendon-driven continuum manipulator. The curved tendon paths deform the body in various shapes. Starting from the classical CRT, they derive the expression of distributed loads applied by tendons on the backbone . The model validation considers straight and non-straight tendon routings, point loads, and distributed external loads . The simulated matched experimental tip positions with a 1.7%L average error. Another experimental work presented a continuum arm with a Nitinol backbone and three tendons turning around the centerline with a constant radial distance (Dehghani and Moosavian, 2013), which extended a previous setup lacking an actuation system (Dehghani and Moosavian, 2011). They modified differential terms of an existing CRT (Rucker and Webster III, 2011), retaining the original accuracy with reduced numerical effort. The model was extended to real-time control using a stable moment-based algorithm to solve the BVP considering the physical characteristics of the rod (Dehghani and Moosavian, 2014a). They also propose a method to reduce the cumulative numerical integration errors that could make the rotation matrices non-orthogonal. Finally, they use the model to develop a Jacobian-based quasi-static tip position controller for a multi-segment continuum arm.
Cosserat rods can simulate the mechanics of several robotic systems, including robotic tails and octopus-inspired arms. The work in (Rone and Ben-Tzvi, 2014) used a Cosserat rod to study a cable-driven multi-section continuum robotic tail envisioned for onboard mobile robots. Prescribed tail configurations or trajectories are analyzed, while the governing equations compute the loading at the tail base transmitted to the mobile robot. This study investigates the impact of trajectory and design factors on the resulting loading profiles, considering factors like mode shape, speed, bending magnitude, and bending plane angle. Key findings include that a non-straight tail reference configuration increases the moment magnitude, short proximal segments provide a larger tip motion for a given bending angle, and higher mass distribution towards the final nodes enhances bending moments and control fidelity. Xu et al. (2017) propose a Cosserat rod model to describe the bending and torsion behavior of a bionic tendon-driven octopus tentacle. Given the location of the four cables, the model maps cable tension to expected bending. Exploiting the derived mapping, they propose an open-loop controller for the robot. Similarly, Niu et al. (2019) derive closed-form equations of a soft robot modeled through the CRT. The equations are validated with a cable-driven soft arm inspired by an octopus tentacle, reaching accurate predictions of the end-effector positions. Chikhaoui et al. (2019) investigated the kinematic modeling of a magnetically extensible tendon actuated continuum robot (METAbot) (Nguyen and Burgner-Kahrs, 2015). The comparison between a PCC beam mechanics approach and a Cosserat rod shows that the latter has higher position accuracy, although computationally heavier. Recently, Janabi-Sharifi et al. (2021) presented a tutorial on the dynamic Cosserat rod modeling of tendon-driven continuum robots. The tutorial explains the core assemblies of tendon-driven continuum robots. Then, it provides a unified approach for the dynamic analysis of tendon-driven continuum robots that bend and twist, including multi-section, concentric tubes, and parallel cooperative robots.
4.3.2 Bend & Twist in Soft Pneumatic Actuators
Actuation-induced torsion is not common in fluidic actuators. However, a Cosserat rod model for a CSM with two asymmetric pneumatic actuators that bend and rotate was introduced (Uppalapati et al., 2018). They optimize the rigidity parameters and the curvatures for quasi-static shapes. The model reproduced the experimental static workspace with a tip load attached. The robot’s dexterity was enhanced integrating an additional soft rotating actuator (Uppalapati and krishnan, 2021). The CRT predicts the deformed robot configuration subjected to external loads, and they experimentally determine elasticity and actuation parameters.
Another study introduced a multi-segment Euler-Bernoulli rod model to predict the bending deformation of micro-tube pneumatic actuators (Ji et al., 2019). They first employ the line-segment model and then extend it to a multi-segment version of Euler-Bernoulli’s rod model for a rapid design, analysis, and optimization of the bending deformation of such actuators. Finally, an analytical model of a generalized PneuNet soft actuator combines the minimum potential energy method and an Euler-Bernoulli rod (Gu et al., 2021). The study compared the rod model to FEM simulations and experimental results, achieving accurate deformed shape predictions, given the geometry and the applied pressures.
4.3.3 Bend & Twist in Tubular Robots
Biomedical applications could leverage CTR and eccentric tube robots that bend and twist. A geometrically exact KRT solved the statics problem for CTR, describing the shape of a collection of concentric pre-curved tubes under general wrench distributions (Rucker et al., 2010a). The model is validated experimentally for an active cannula with three concentric, telescopic, pre-curved Nitinol tubes in several configurations with varying tip loads. The results establish the basis for model-based position and force control of active cannulas and their use as force sensors under external loads. Also, they tackle the integration of unmodelled effects such as friction, elongation, and transverse shear (Rucker et al., 2010b). Subsequently, Till et al. (2020) present the first dynamic model for CTR. The model describes the coupled inertial dynamics of pre-curved concentric tubes considering the dynamic effects of linear and rotational tube inertia, material damping, Coulomb and viscous friction, and the inertia of a rigid body at the robot tip. The experimental validation measures the dynamic behavior during tissue grasping and snapping.
The structure of the ETR extends the CTR design. A kinematic model of a two-channel ETR with two continuum arms delivered through the channels is presented (Wang et al., 2019). Pre-curved superelastic tubes lining the arm’s working channels produce sheath deflection. Cosseat rods model the pre-curved tubes and the central backbone, achieving promising tip pose predictions for pure rotation actuation. Similarly, a continuum ETR sheath for MIS such as neuroendoscopy, requiring delivery of multiple robotic arms, was presented (Wang et al., 2021a). Super-elastic tubes lining the arm channels actuate the sheath, while distributed disks are attached to the backbone. The tube kinematics and the robot backbone are modeled using the CRT. Three experiments with a two-arm sheath validate the tip position accuracy. Another work proposes a kinematic Cosserat rod model of a snake-like continuum robot for MIS (Wang et al., 2020). The robot geometry comprises three identical super-elastic Nitinol tubes held eccentrically by a surrounding sheath. Assuming sheat inextensibility, they show that different actuations virtually yield the same robot shape and tip pose. Simulations and experiments validate the model for a single-stage ETR.
Safety is crucial in surgical procedures. A method to compute compliance of continuum robots, modeled as quasi-static Cosserat rods under diverse loads, was devised (Smoljkic et al., 2014). The robot compliance is described by ODEs and integrated with the forward kinematics. The experiments involve a non-actuated flexible rod composed of a super-elastic Nitinol tube surrounding a braided Pebax tube. The rod mounted on a force sensor with position sensors embedded in the tip is subject to external tip forces. A least square fitting of the measured tip trajectory versus the direct kinematics identifies material parameters and tube configuration. Finally, they compared the velocities obtained through position differentiation against those obtained multiplying the differentiated forces and the computed compliance matrices.
4.3.4 Bend & Twist with Smart Actuators
The KRT modeled TCP emphasizing critical factors like coil load and twist and incorporating them into the actuation equations for TCP muscles (Wu and Zheng, 2020). The model captures the quantitative relationship between tensile actuation and fabrication load, serving as a predictor for tensile stroke. Also, it was validated through TCP muscles made of silver-coated Nylon fiber under different fabrication loads, measuring recovery stress and maximum stroke during heating. Thermo-mechanical tension tests conducted at various room temperatures identified the elastic modulus of the fibers. To achieve variable stiffness in continuum robots, Wang et al. (2022b) propose to use winding-styled ropes inspired by snakes’ behavior. Controlling the temperature of an SMA spring increased their stiffness up to . The experimental validation of a kinetic-static Cosserat rod model obtained an error below .
4.3.5 Bend & Twist in Parallel Manipulators
4.3.6 Bend & Twist with Buckling Instability
Buckling instability is an important factor to consider in soft robots. With applications in soft robotics, Sipos and Várkonyi (2020) explores the interplay between spatial buckling and arbitrary intrinsic curvature on a simulated clamped Cosserat rod with a vertical tip load. The aim is to tune the control parameters to find the highest load for a given free-end position or the maximum reaching of the cantilever in a direction for a given load. The numerical investigation of the spatial instability of in-plane configurations of soft elastic cantilevers subject to an endpoint load demonstrates the importance of out-of-plane instability. Moreover, the capability of controlled intrinsic curvatures to stabilize the rod in configurations of high reaching is shown. Significant improvements occur for flattish cross-sections where the rigidities associated with out-of-plane bending and torsion are larger than the in-plane stiffness. At the same time, arms with high and narrow cross-sections are slightly more performant. Buckling instability can be triggered. Forghani et al. (2021) analyze the inverse dynamics and control of a bacteria-inspired uniflagellar robot in a fluid medium at a low Reynolds number. The robot, mimicking the locomotion of flagellated bacteria, features a flexible helical filament (flagellum) connected to a spherical head. A Kirchhoff elastic rod describes the flagellum, which rotates about the head at a controlled angular velocity and generates propulsive forces for the forward motion. The results demonstrate that by triggering the buckling instability of the flagellum, the robot follows a prescribed path in space. The control scheme only involves the flagellum angular velocity, and a deep neural network captures the dynamics predicting the robot’s trajectory. However, higher maneuverability tasks like drug delivery require multiple flagella, advanced controllers, and control inputs.
4.3.7 Bend & Twist in Insect-like Wings
The applicability of rod theories goes beyond manipulators. A study presented a hybrid dynamic Cosserat rod model of a 3D aerial insect-like robot with two soft wings (Belkhiri et al., 2012). In particular, Cosserat rods described the leading edge of the wings, which bend and twist based on continuous Newton-Euler dynamics. The wings, connected to the insect thorax through revolute joints, are treated as rigid bodies. The hybrid model combines discrete and continuous models considering the curvature deformations and external aerodynamic forces. The model replicates the kinematics and internal deformations of the wings and the rigid net insect motion produced by a flapping flight.
Subclass | References | Rod Model | Contribution |
Cable | Rucker and Webster III (2011) | CRT | General dynamic model of a continuum manipulator with curvilinear actuators. |
@@ 4.3.7 Bend & Twist in Insect-like Wings
Pneumatic | Uppalapati and krishnan (2021) | CRT | Design and modeling of a CSM with parallel asymmetric fiber-reinforced elastomers. |
@@
Tubular | Till et al. (2020) | CRT | Dynamic model of CTR validated for tissue grasping and snapping. |
Wang et al. (2021a) | CRT | Modeling of a superelastic ETR sheath for biomedical procedures. | |
@@
Smart | Wu and Zheng (2020) | KRT | TCP muscles incorporating coil load and twist into actuation model. |

### v2

@@ 6.2 All Deformations
With the advancement of soft robotics, the research community started exploring the use of continuum and soft robots in unstructured environments. In particular, when interacting with irregularly shaped objects, shear deformation cannot be neglected. Additionally, advancements in actuation technologies necessitate the consideration of stretch and torsion modes. Within this context, the CRT emerged as a fundamental tool for modeling the deformation of continuum and soft robots. A recent contribution by Tuummers2023Cosserat further consolidates CRT as a standard framework, by rigorously deriving both Newtonian and Lagrangian formulations for multi-segment tendon-driven continuum robots and demonstrating their equivalence, with practical guidelines for numerical implementation.
As discussed in Sec. 4, the CRT describes the motion of an elastic rod with an infinite number of DoFs. Discretization techniques play a key role in numerically solving the governing equations of Cosserat rods. Several methods were developed to enhance numerical robustness and computational efficiency. For example, renda2018discrete introduced the piecewise constant strain (PCS) method, which discretizes a continuous Cosserat rod into segments with constant strain. This method extends the PCC approach (webster2010design) by incorporating shear and torsion. In parallel, till2019real proposed a real-time discretization technique that transforms the PDE of a Cosserat rod into a boundary value problem (BVP) by discretizing time derivatives, significantly improving computational performance.
Incorporating external forces such as buoyancy, drag, and added mass is essential for accurately modeling the interactions between soft robots and their surroundings. In this context, renda20123d; renda2014dynamic analyzed the static and dynamic behavior of a silicone arm in a dense medium, capturing its coupled tendon conditions and external loads. The model was further developed in a unified Cosserat rod framework applied to underwater robots (renda2018unified). By accounting for nonlinearities and hydrodynamic effects, this approach enabled the evaluation of soft robot designs under submerged conditions.
For pneumatic actuators, renda2017screw extended the PCS method by integrating Screw Theory, allowing flexible routing of fluidic and tendon actuators along arbitrary paths. In PCM, armanini2021discrete adapted the PCS formulation to closed-chain soft robotic systems, facilitating the modeling and design optimization of structures like the Fin-Ray finger.
Material nonlinearities, such as hysteresis, can significantly affect the performance of soft robots. mishra2022fractional tackled the challenge of large deformations in pneumatic soft robots by proposing a fractional-order Bouc-Wen model within a Cosserat rod framework. By incorporating nonlinear material effects, the model demonstrated superior accuracy in capturing the behavior of soft actuators compared to traditional approaches.
Extensive validation of CRT models is essential for ensuring their predictive accuracy. Cable-driven soft robots were widely studied and experimentally validated. For example, renda20123d validated a static Cosserat model of a tendon-driven soft arm using FEM simulations and experimental comparisons. The dynamic extension of this model (renda2014dynamic) was experimentally verified by replicating characteristic octopus movements. Later, renda2018unified demonstrated the effectiveness of their unified dynamic framework using an underwater robotic system. Further validation of discretization-based approaches was performed in renda2018discrete and till2019real, confirming the numerical robustness of their methods. Conversely, butler2019continuum investigated deflection and stiffness in robots subject to constant axial tendon displacement and environmental loads, comparing parallel and converging actuator paths. They validated the Cosserat rod model on a CSM subject to tip loads in the kg range, providing actuations to reach bending angles up to . The simulation results aligned well with the experimental data. However, as the authors noted, limitations emerged for larger curvatures due to the increased role of friction, which was not included in the model. For pneumatic soft robots, trivedi2007geometrically; trivedi2008geometrically developed geometrically exact CRT models and validated them on fiber-reinforced pneumatic manipulators, achieving significantly higher accuracy than constant curvature models. The screw-theoretic PCS formulation (renda2017screw) was validated on the STIFF-FLOP arm, while Roshanfar2021towards tested their pneumatic Cosserat rod model experimentally, reporting an tip position error. Also, FEM was extensively used to compare and validate CRT models. Indeed, grazioso2018analytic derived analytical solutions for quasi-static Cosserat rods and verified them against FEM simulations under external loads. Similarly, berthold2021preliminary conducted a comparative study between FEM and CRT models for pneumatic actuators, proposing a parameter identification process. Further refinement was made by wiese2022towards, who obtained a static Cosserat model from detailed FEM simulations, validating it experimentally across various loading conditions. For parallel continuum robots, till2017elastic investigated elastic stability using Cosserat rod models and experimentally validated their stability conditions. In parallel, black2017parallel introduced a kinetostatic CRT framework for PCM, achieving high accuracy in force sensing and manipulability analysis. Then, lilge2022kinetostatic extended this approach, developing a comprehensive CRT-based framework for tendon-driven PCM, experimentally achieving a median pose accuracy of .
Cosserat rod models found widespread applications in bioinspired robotics and biomechanical studies. boyer2017locomotion proposed a unified CRT-based framework for modeling the locomotion of bioinspired robots with soft appendages. They analyzed flapping flight and passive swimming, comparing two different modeling approaches based on Newton-Euler dynamics. For musculoskeletal modeling, zhang2019modeling implemented Cosserat rod simulations in PyElastica (naughton2021elastica), assembling heterogeneous structures to replicate musculoskeletal architectures (e.g., snakes, bird wings). Their work demonstrated the effectiveness of CRT in bioengineering design, simulating complex musculoskeletal interactions in biological and robotic applications. boyer2022statics explored the intersection of CRT and optimal control theory, revealing fundamental insights into the singularity of dynamic continuum robot simulations. By reformulating the problem as a minimization task, their approach provided a novel perspective on the statics and dynamics of continuum robots, validated through simulations of bioinspired continuum swimmers.
Finally, to find optimal bases for the GVS approach, alkayas2024soft introduced a data-driven reduction method based on Proper Orthogonal Decomposition (POD). Their approach begins by creating a high-fidelity GVS digital twin from experimental data, which is then simulated to generate a comprehensive strain dataset. By applying Singular Value Decomposition (SVD) to this data and truncating the least significant singular values, the model’s dimensionality is reduced while preserving its accuracy. This process yields an optimal linear basis of the form . Building on this work, a subsequent study extended the search to the nonlinear case (i.e., ), employing autoencoders to identify a more expressive and compact representation (alkayas2025structure).
@@ 6.3 Bend & Twist
Bending and twisting are the fundamental deformation modes of CSM, enabling them to achieve complex 3D shapes and dexterous movements. Modeling these modes is particularly important in applications where interaction with the environment is not limited to planar motions but requires full spatial maneuverability. In many cases, incorporating twisting alongside bending provides the best trade-off between accuracy and computational efficiency. This allows continuum robots to follow arbitrary 3D curves in space, which is essential for tasks such as wrapping around an object, executing dexterous maneuvers, or navigating constrained environments. Biomedical applications especially benefit from bending-twisting modes, as in concentric tube robots for minimally invasive surgery (MIS) (rucker2010model; rucker2010geometrically; till2020dynamic) and eccentric tube robots for multi-arm sheaths (wang2019steering; wang2020eccentric; wang2021eccentric). Safety-critical domains also utilize bending-twisting compliance analysis to predict deformations under external loads (smoljkic2014compliance).
Twisting deformations are usually induced by particular cable or tendon routing. Tendon-driven continuum robots employ non-straight tendon paths that wrap around the backbone to generate both bending and torsional modes. Early work by rucker2011statics established Cosserat rod formulations where curved tendon paths produce distributed loads along the backbone. Similarly, dehghani2013modeling developed a Nitinol-backbone continuum arm with three tendons spiraling around the centerline, extending their previous untwisted setup (dehghani2011mod_cont). They further refined this approach for real-time control using a stable moment-based algorithm (dehghani2014compact). Twisting was also demonstrated in robotic tails (rone2014continuum), octopus-inspired manipulators (xu2017modelling; niu2019closed), tendon-actuated extensible robots (chikhaoui2019comparison), and multi-section tendon robots (janabi2021cosserat). In fluidic actuation, torsional effects are less common but can be induced through asymmetric chambers or fiber reinforcement. For example, uppalapati2018parameter modeled a manipulator with two asymmetric pneumatic actuators capable of bending and rotating, which were later extended with an additional rotating actuator (uppalapati2021design). Other approaches include models based on Euler-Bernoulli for micro-tube pneumatics (ji2019rapid) and generalized PneuNet structures, combining bending and twisting predictions validated against FEM (gu2021analytical).
Cosserat rods were extensively applied to tendon-driven systems, while Euler-Bernoulli or energy-based methods are often preferred for pneumatic networks due to their simpler kinematics. Examples include asymmetric pneumatic actuators capable of torsion (uppalapati2018parameter; uppalapati2021design), micro-tube pneumatics modeled with multi-segment Euler–Bernoulli rods (ji2019rapid), and PneuNet actuators modeled analytically with energy minimization (gu2021analytical). These models capture how chamber geometry, reinforcement, and pressurization induce coupled bending and twisting modes. However, the majority of bending-twisting continuum robots are tendon-driven, concentric-tube, or eccentric-tube systems. These architectures require a geometrically exact description of distributed loads and torsional effects, which is naturally expressed through the CRT.
Concerning CTR, renda2021sliding adapted the piecewise variable-strain (mathew2025reduced) to simulate multi-section CTR by including the tubes’ sliding motion. Rotation motions of the tubes are included as generalized coordinates rather than boundary kinematic conditions. Notably, this procedure led to a minimal set of closed-form algebraic equations solvable for both the shape variables, the actuation forces, and the torques. The resulting equations also facilitate control, design optimization, and stability assessment. Unlike beam simplifications, Cosserat rods account for coupling between bending and torsion and handle spatially varying loads from tendon paths, external wrenches, and contact. This explains why most bending-twisting works in the literature adopt the Cosserat rod formalism (rucker2011statics; dehghani2013modeling; dehghani2014compact; rone2014continuum; xu2017modelling; niu2019closed; chikhaoui2019comparison; janabi2021cosserat; wang2019steering; wang2020eccentric; wang2021eccentric).
Buckling instability is another important nonlinear phenomenon to consider. When modeling thermally activated twisted and coiled polymer (TCP) muscles, wu2020modeling employed KRT, since the small-strain assumptions and direct link to coil load-twist relations simplify the actuation model. Similarly, Kirchhoff rods were used in micrometer-scale parallel continuum robots (mauze2021micrometer) and bacteria-inspired flagellated swimmers exploiting buckling instabilities for propulsion (forghani2021control). Conversely, Cosserat rods were applied to study global buckling in soft robotic arms (sipos2020longest), where large deflections and intrinsic curvature control are essential. Thus, Kirchhoff rods often appear in actuation-centric or microscale studies where analytical tractability is crucial, while Cosserat rods dominate in large-deformation, shape-control analyses.
Finally, workspace analysis requires accurate models of bending and twisting under actuation and loading. Tendon-driven Cosserat models enable forward kinematics and Jacobian-based control for workspace prediction (dehghani2014compact). For pneumatic robots, static workspace matching was validated experimentally with asymmetric actuators modeled by Cosserat rods (uppalapati2018parameter; uppalapati2021design). PneuNet and micro-tube actuator models (ji2019rapid; gu2021analytical) allow for workspace design and optimization by linking actuator geometry and pressure inputs to achievable curves. In biomedical contexts, concentric and eccentric tube models predict reachable poses in confined environments (rucker2010model; wang2021eccentric).
| Deformation Class | Subclass | Reference | Contribution | Rod Model |
| Bend | Cable | jones2009three_dim | Static model for real-time shape estimation of CSM. | CRT |
| dehghani2014dynamics | Dynamic model of a continuum finger for planar grasping. | CRT | ||
| gao2016mechanical | Model of notched continuum manipulator for MIS. | CRT | ||
| Pneumatic | de2017development | Derivation of constitutive law of a planar bending actuator. | EBRT | |
| sadati2019reduced | ROM and rod models with absolute and relative states validated on STIFF-FLOP. | EBRT | ||
| Smart | daily2017dynamical | Dynamic model inspired by caterpillar motion to analyse ondulation mechanics. | EBRT | |
| Pawlowski2018modeling | Modeling the coupling between CSM body and thermal actuation. | CRT | ||
| Parallel | li2018shape | Shape modeling of a PCM made of soft panels. | EBRT | |
| Tubular | pattanshetti2019kinematic | Estimation of stiffness variation on a monolithic tube for MIS. | CRT | |
| All Deformations | Cable | rucker2011statics | Derived exact models for the forward kinematics, statics, and dynamics of CSM with general tendon routing. | CRT |
| renda2014dynamic | Studied the dynamic interaction of an octopus-inspired CSM with a dense medium. | |||
| butler2019continuum | Predicted deflection and stiffness in CSM subject to constant actuation and external loads. | |||
| Pneumatic | trivedi2008geometrically | Model validation of CSM with fiber-reinforced actuators varying base orientation. | ||
| Parallel | till2017elastic | Studied and experimentally validated the elastic stability of PCM. | ||
| armanini2021discrete | Modeled closed-chain soft PCM with validation on a soft-rigid finger. | |||
| lilge2022kinetostatic | Studied reachability, singularities, manipulability, and compliance of PCM. | |||
| Bioinspired | zhang2019modeling | Simulates musculoskeletal systems assembling heterogeneous active/passive rods. | ||
| boyer2022statics | Studied the link between CRT and optimal control simulating locomotors and swimmers. | |||
| Bend & Twist | Cable | janabi2021cosserat | Tutorial on the dynamic modeling of tendon-driven continuum robots that bend and twist. | CRT |
@@
| Pneumatic | uppalapati2021design | Design and modeling of a CSM with parallel asymmetric fiber-reinforced elastomers. | CRT | |
@@
| Tubular | till2020dynamic | Dynamic model of CTR validated for tissue grasping and snapping. | CRT | |
| wang2021eccentric | Modeling of a superelastic ETR sheath for biomedical procedures. | CRT | ||
| renda2021sliding | Adapted the piecewise variable strain to CTR including the tubes’ sliding motion. | CRT | ||
@@
| Smart | wu2020modeling | TCP muscles incorporating coil load and twist into actuation model. | KRT | |